OpenStudy (anonymous):

PLEASE HELP!!! Simplify 20 sin (2x) cos (2x)

zepdrix (zepdrix):

Sine Double Angle again, \(\large\rm 2\sin(\color{orangered}{\theta})\cos(\color{orangered}{\theta})=\sin(2\cdot\color{orangered}{\theta})\)

zepdrix (zepdrix):

But we don't want a 2 on the left side of the equation, we want a 20, right? Let's multiply both sides by 10,\[\large\rm 20\sin(\color{orangered}{\theta})\cos(\color{orangered}{\theta})=10\sin(2\cdot\color{orangered}{\theta})\]

zepdrix (zepdrix):

You finish from there :)

OpenStudy (anonymous):

Thank you!!